# Fisher-Rao Gradient Flow: Geodesic Convexity and Functional Inequalities

### v1

@@ 3.1. Preliminaries on convex analysis
(3.1) $h(\lambda\theta_{0}+(1-\lambda)\theta_{1})\leq\lambda h(\theta_{0})+(1-\lambda)h(\theta_{1}).$
@@
(3.2) $h(\theta_{1})\geq h(\theta_{0})+\langle\nabla_{\theta}h(\theta_{0}),\theta_{1}-\theta_{0}\rangle,$
@@
(3.3) $h(\theta_{1})\geq h(\theta_{0})+\langle\nabla_{\theta}h(\theta_{0}),\theta_{1}-\theta_{0}\rangle+\frac{\alpha}{2}|\theta_{1}-\theta_{0}|^{2}.$
For a gradient flow in the Euclidean space, denoted by
(3.4) $\frac{{\rm d}\theta_{t}}{{\rm d}t}=-\nabla_{\theta}h(\theta_{t}),$
the evolution of the function values satisfy
(3.5) $\frac{{\rm d}}{{\rm d}t}h(\theta_{t})=-|\nabla_{\theta}h(\theta_{t})|^{2}.$
Equation 3.3 indicates that $h$ has a unique global minimizer, denoted by $\theta^{*}$. Using the strong convexity of $h$ we have
(3.6) $h(\theta)-h({\theta^{*}})\leq\langle\nabla_{\theta}h(\theta),\theta-{\theta^{*}}\rangle-\frac{\alpha}{2}|{\theta^{*}}-\theta|^{2}\leq\frac{1}{2\alpha}|\nabla_{\theta}h(\theta)|^{2},$
@@
(3.7) $|\nabla_{\theta}h(\theta)|^{2}\geq 2\alpha(h(\theta)-h({\theta^{*}})),$
which is referred to as the gradient dominance condition or the Polyak-Łojasiewicz inequality [73]. With this inequality, the gradient flow converges exponentially fast, as $\frac{{\rm d}}{{\rm d}t}h(\theta_{t})\leq-2\alpha(h(\theta_{t})-h({\theta^{*}}))$ leading to $h(\theta_{t})-h({\theta^{*}})=\mathcal{O}(\exp(-2\alpha t))$. In a summary, the strong convexity in the Euclidean space leads to gradient dominance condition, which then implies the exponential convergence of the gradient flow.
@@
(3.8) $\mathcal{D}^{2}(p_{0},p_{1})=\inf_{p_{t}\in\mathcal{M},v_{t}\in T_{p}\mathcal{M}}\{\int_{0}^{1}g_{p_{t}}(v_{t},v_{t}){\rm d}t:\ \frac{{\rm d}}{{\rm d}t}p_{t}=v_{t},p_{t=0}=p_{0},p_{t=1}=p_{1}\},$
where $\mathcal{D}$ is called the geodesic distance between $p_{0}$ and $p_{1}$. Such a geodesics satisfies $g_{p_{t}}(v_{t},v_{t})=\mathcal{D}^{2}(p_{0},p_{1})$ so is referred to as the constant-speed geodesics.
@@
(3.9) $h(p_{\lambda})\leq\lambda h(p_{0})+(1-\lambda)h(p_{1}),$
@@
(3.10) $h(p_{1})\geq h(p_{0})+g_{p_{0}}(\text{grad}~{}h(p_{0}),v_{0})+\frac{\alpha}{2}\mathcal{D}^{2}(p_{0},p_{1}),$
@@
(3.11) $g_{p_{0}}(\textrm{Hess}\,h(p_{0})v_{0},v_{0})=\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\Big{|}_{\rm geod}h(p_{0})$
@@
Similarly as the Euclidean case, $\alpha$-strongly geodesic convexity implies the gradient dominance condition:
(3.12) $g_{p}(\text{grad}\,h(p),\text{grad}\,h(p))\geq 2\alpha(h(p)-h(p^{*})),$
where $p^{*}$ is the global minimizer of $h$. Again, using this inequality leads to the exponential convergence of gradient flows of $h$ on the manifold $\mathcal{M}$.
3.2. Hessian operator in the Fisher-Rao geometry
@@
(3.13) $g_{\rho}^{\mathrm{FR}}(\sigma_{1},\sigma_{2})=\int\frac{\sigma_{1}\sigma_{2}}{\rho}\mathrm{d}\theta,\text{ for }\sigma_{1},\sigma_{2}\in T_{\rho}\mathcal{P}.$
The geodesic distance between $\rho_{1},\rho_{1}$, referred to as the Fisher-Rao distance, satisfies
(3.14a) $\mathcal{D}_{\rm FR}^{2}(\rho_{0},\rho_{1})$ $=\inf_{\rho_{t},\sigma_{t}}\Bigl{\{}\int_{0}^{1}\int\frac{|\sigma_{t}|^{2}}{\rho_{t}}\mathrm{d}\theta\mathrm{d}t:\partial_{t}\rho_{t}=\sigma_{t}\Bigr{\}}$
(3.14b) $=\inf_{\rho_{t},\psi_{t}}\Bigl{\{}\int_{0}^{1}\int\rho_{t}(\psi_{t}-\mathbb{E}_{\rho_{t}}[\psi_{t}])^{2}\mathrm{d}\theta\mathrm{d}t:\partial_{t}\rho_{t}=\rho_{t}(\psi_{t}-\mathbb{E}_{\rho_{t}}[\psi_{t}])\Bigr{\}}.$
Remark 3.1.
The Fisher-Rao geodesic distance attains an explicit formula, equivalent to (up to a constant scaling) the spherical Hellinger distance:
$\mathcal{D}_{\rm FR}^{2}(\rho_{0},\rho_{1})=4\operatorname{arccos}^{2}\left(\int\sqrt{\rho_{0}}\sqrt{\rho_{1}}{\rm d}\theta\right).$
The constant-speed geodesics is
$\rho_{t}(\theta)=\frac{\left(t\sqrt{\rho_{0}(\theta)}+(1-t)\sqrt{\rho_{1}(\theta)}\right)^{2}}{(1-t)^{2}+t^{2}+2t(1-t)\int\sqrt{\rho_{0}(\theta^{\prime})\rho_{1}(\theta^{\prime})}{\rm d}\theta^{\prime}}.$
For more discussions, see [40, 53, 60]. In view of this relation, Fisher-Rao gradient flows are sometimes referred to as spherical Hellinger gradient flows in the literature [56, 60].
$\lozenge$
Remark 3.2.
If we do not restrict the distributions $\rho_{t}$ to be on the probability space and we allow them to have any positive mass, then by using the relation
$\frac{|\sigma_{t}|^{2}}{\rho_{t}}=\frac{|\partial_{t}{\rho}_{t}|^{2}}{\rho_{t}}=4\Bigl{|}\frac{\mathrm{d}}{\mathrm{d}t}\sqrt{\rho_{t}}\Bigr{|}^{2}$
and the Cauchy-Schwarz inequality, we will get that the optimal objective value in (3.14) is $4\int|\sqrt{\rho_{0}}-\sqrt{\rho_{1}}|^{2}\mathrm{d}\theta.$ This is (up to a constant scaling) the Hellinger distance [37]. $\lozenge$
Using a formal Lagrangian duality calculation, we can derive the optimality condition of the variational problem (3.14) as a coupled system of differential equations (see also [83]):
(3.15a) $\partial_{t}\rho_{t}$ $=\rho_{t}(\psi_{t}-\mathbb{E}_{\rho_{t}}[\psi_{t}]),$
(3.15b) $\partial_{t}\psi_{t}$ $=-\frac{1}{2}(\psi_{t}-\mathbb{E}_{\rho_{t}}[\psi_{t}])^{2},$
@@
(3.16a) $g_{\rho}^{\rm FR}(\textrm{grad}_{\rm FR}\mathcal{E}(\rho),\sigma)=$ $\frac{\mathrm{d}}{\mathrm{d}t}\Big{|}_{\rm geod}\mathcal{E}(\rho)=\int\frac{\delta\mathcal{E}}{\delta\rho}\rho(\psi-\mathbb{E}_{\rho}[\psi])\mathrm{d}\theta,$
$g_{\rho}^{\rm FR}(\textrm{Hess}_{\rm FR}\mathcal{E}(\rho)\sigma,\sigma)=$ $\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\Big{|}_{\rm geod}\mathcal{E}(\rho)=\frac{\mathrm{d}}{\mathrm{d}t}\Big{|}_{\rm geod}\int\frac{\delta\mathcal{E}}{\delta\rho}\rho(\psi-\mathbb{E}_{\rho}[\psi])\mathrm{d}\theta$
(3.16b) $=$ $\int\frac{\delta^{2}\mathcal{E}}{\delta\rho^{2}}\rho^{2}(\psi-\mathbb{E}_{\rho}[\psi])^{2}+\frac{1}{2}\rho(\psi-\mathbb{E}_{\rho}[\psi])^{2}\Bigl{(}\frac{\delta\mathcal{E}}{\delta\rho}-\mathbb{E}_{\rho}[\frac{\delta\mathcal{E}}{\delta\rho}]\Bigr{)}\mathrm{d}\theta.$
3.3. Geodesic convexity in the Fisher-Rao geometry
For the $f$-divergences (2.1), the corresponding gradient and Hessian operators are
(3.17a) $g_{\rho}^{\rm FR}(\textrm{grad}_{\rm FR}D_{f}[\rho\|{\rho^{*}}],\sigma)=$ $\int f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}\rho(\psi-\mathbb{E}_{\rho}[\psi])\mathrm{d}\theta,$
(3.17b) $g_{\rho}^{\rm FR}(\textrm{Hess}_{\rm FR}D_{f}[\rho\|{\rho^{*}}]\sigma,\sigma)=$ $\int f^{\prime\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}\frac{\rho^{2}(\psi-\mathbb{E}_{\rho}[\psi])^{2}}{{\rho^{*}}}$
$+$ $\frac{1}{2}\rho(\psi-\mathbb{E}_{\rho}[\psi])^{2}\Bigl{(}f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}-\mathbb{E}_{\rho}[f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}]\Bigr{)}\mathrm{d}\theta,$
where $\sigma=\rho(\psi-\mathbb{E}_{\rho}[\psi])$.
@@ 3.3. Geodesic convexity in the Fisher-Rao geometry
Proof.
The KL-divergence corresponds to the $f$-divergence with $f(x)=x\log x$. In such case, the Hessian formula (3.17b) has the form
(3.18) $\begin{split}g_{\rho}^{\rm FR}(\textrm{Hess}_{\rm FR}{\rm KL}[\rho\|{\rho^{*}}]\sigma,\sigma)&=\int\rho(\psi-\mathbb{E}_{\rho}[\psi])^{2}\\ &+\frac{1}{2}\rho(\psi-\mathbb{E}_{\rho}[\psi])^{2}\Bigl{(}\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr{)}\mathrm{d}\theta.\end{split}$
Thus, whether the Hessian is positive semi-definite or not depends on the sign of $1+\frac{1}{2}(\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}])$. The KL divergence will be geodesically convex when this term is non-negative. A simple sufficient condition is $1/e\leq{\rho}/{{\rho^{*}}}\leq e$.
However, in general, this term does not have a definite sign, and the KL divergence is not geodesically convex. To show that the Hessian is not positive semi-definite, we set a specific $\psi=\log\frac{\rho}{{\rho^{*}}}$, then (3.18) becomes
(3.19) $\mathcal{H}(\rho,{\rho^{*}})=\int\rho\Bigl{(}\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr{)}^{2}+\frac{1}{2}\rho\Bigl{(}\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr{)}^{3}\mathrm{d}\theta.$
We construct the following two examples to explain why the above term is not non-negative in general.
We first consider 1D Gaussian densities $\rho(\theta)=\mathcal{N}(\theta;\mu,\sigma^{2})$ and ${\rho^{*}}(\theta)=\mathcal{N}(\theta;0,1)$. We have that
(3.20) $\begin{split}&\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]=\frac{\sigma^{2}}{2}+\frac{\mu^{2}}{2}-\log\sigma-\frac{1}{2},\\ &\int\rho\Bigl{(}\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr{)}^{2}\mathrm{d}\theta=\frac{1}{2}(\sigma^{2}-1)^{2}+\sigma^{2}\mu^{2},\\ &\int\rho\Bigl{(}\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr{)}^{3}\mathrm{d}\theta=(\sigma^{2}-1)^{3}+3(\sigma^{2}-1)\sigma^{2}\mu^{2}.\end{split}$
Plugging (3.20) into (3.19) leads to that
@@
Therefore, when
$\sigma^{2}<\frac{1}{3}\qquad\textrm{and}\qquad\mu^{2}>\frac{(\sigma^{2}-1)^{2}}{1-3\sigma^{2}},$
we have that $\mathcal{H}(\rho,{\rho^{*}})<0$.
Next, we discuss a general example beyond the Gaussian case. We will first give a counterexample where the density is not continuous. We then discuss the construction of a counterexample with a smooth density, by convolving the previous example with a smooth bump function.
@@
(3.21) $\rho_{r}:=\int_{S_{r}}{\rho^{*}}(\theta)\mathrm{d}\theta,$
@@
(3.22) $\rho(\theta)={\rho^{*}}(\theta)(x_{1}\mathbbm{1}_{S^{c}_{r}}(\theta)+x_{2}\mathbbm{1}_{S_{r}}(\theta))$
where $\mathbbm{1}_{A}$ is the indicator function of a set $A$; $S_{r}$ denotes the sphere with radius $r$. For $\rho(\theta)$ to be a probability density, it is necessary that
(3.23) $x_{1}\Bigl{(}1-\int_{S_{r}}{\rho^{*}}(\theta)\mathrm{d}\theta\Bigr{)}+x_{2}\int_{S_{r}}{\rho^{*}}(\theta)\mathrm{d}\theta=x_{1}(1-\rho_{r})+x_{2}\rho_{r}=1.$
To see why this is the case, note that we can vary $r\in(0,+\infty)$ such that $\rho_{r}$ in (3.21) can take any value in $(0,1)$. Thus, for any choice of $x_{1}>1>x_{2}>0$, we can take some $r\in(0,+\infty)$ such that
$\rho_{r}=\frac{x_{1}-1}{x_{1}-x_{2}}\in(0,1),$
which leads to (3.23).
We can compute KL$[\rho\|{\rho^{*}}]$ and $\mathcal{H}(\rho,{\rho^{*}})$ denoted in (3.19), as follows:
@@
(3.24a) ${\rm KL}[\rho\|{\rho^{*}}]\leq$ $\frac{x_{1}-x_{1}x_{2}}{x_{1}-x_{2}}\log x_{1}\leq\log x_{1}=\epsilon,$
(3.24b) $\mathcal{H}(\rho,{\rho^{*}})=$ $\rho_{r}(1-\rho_{r})x_{1}x_{2}\Bigl{(}\log\frac{x_{1}}{x_{2}}\Bigr{)}^{2}\Bigl{(}1-\frac{x_{1}+x_{2}-2x_{1}x_{2}}{2(x_{1}-x_{2})}(M+\epsilon)\Bigr{)}$

### v2

@@ 3.1 Preliminaries on convex analysis
(28) $h\bigl(\lambda\theta_{0}+(1-\lambda)\theta_{1}\bigr)\leq\lambda h(\theta_{0})+(1-\lambda)h(\theta_{1}).$
@@
(29) $h(\theta_{1})\geq h(\theta_{0})+\langle\nabla_{\theta}h(\theta_{0}),\theta_{1}-\theta_{0}\rangle,$
@@
(30) $h(\theta_{1})\geq h(\theta_{0})+\langle\nabla_{\theta}h(\theta_{0}),\theta_{1}-\theta_{0}\rangle+\frac{\alpha}{2}|\theta_{1}-\theta_{0}|^{2}.$
For a gradient flow in the Euclidean space, denoted by $\frac{{\rm d}\theta_{t}}{{\rm d}t}=-\nabla_{\theta}h(\theta_{t})$, the evolution of the function values satisfy $\frac{{\rm d}}{{\rm d}t}h(\theta_{t})=-|\nabla_{\theta}h(\theta_{t})|^{2}$. Equation 30 indicates that $h$ has a unique global minimizer, denoted by $\theta^{*}$. Using the strong convexity of $h$ we have
(31) $h(\theta)-h({\theta^{*}})\leq\langle\nabla_{\theta}h(\theta),\theta-{\theta^{*}}\rangle-\frac{\alpha}{2}|{\theta^{*}}-\theta|^{2}\leq\frac{1}{2\alpha}|\nabla_{\theta}h(\theta)|^{2},$
@@
(32) $|\nabla_{\theta}h(\theta)|^{2}\geq 2\alpha\bigl(h(\theta)-h({\theta^{*}})\bigr),$
which is referred to as the gradient dominance condition or the Polyak-Łojasiewicz inequality [polyak1963gradient]. With this inequality, the gradient flow converges exponentially fast, as $\frac{{\rm d}}{{\rm d}t}h(\theta_{t})\leq-2\alpha\bigl(h(\theta_{t})-h({\theta^{*}})\bigr)$ leading to $h(\theta_{t})-h({\theta^{*}})=\mathcal{O}\bigl(\exp(-2\alpha t)\bigr)$. In summary, the strong convexity in the Euclidean space leads to gradient dominance condition, which then implies the exponential convergence of the gradient flow.
@@
(33) $\mathcal{D}^{2}(p_{0},p_{1})=\inf_{p_{t}\in\mathcal{M},v_{t}\in T_{p}\mathcal{M}}\{\int_{0}^{1}g_{p_{t}}(v_{t},v_{t}){\rm d}t:\ \frac{{\rm d}}{{\rm d}t}p_{t}=v_{t},p_{t=0}=p_{0},p_{t=1}=p_{1}\},$
where $\mathcal{D}$ is called the geodesic distance between $p_{0}$ and $p_{1}$, and $v_{t}$ denotes velocity. Such a geodesics satisfies $g_{p_{t}}(v_{t},v_{t})=\mathcal{D}^{2}(p_{0},p_{1})$ so is referred to as the constant-speed geodesics.
@@
(34) $h(p_{\lambda})\leq\lambda h(p_{0})+(1-\lambda)h(p_{1}),$
@@
(35) $h(p_{1})\geq h(p_{0})+g_{p_{0}}\bigl(\text{grad}~h(p_{0}),v_{0}\bigr)+\frac{\alpha}{2}\mathcal{D}^{2}(p_{0},p_{1}),$
@@
(36) $g_{p_{0}}\bigl(\textrm{Hess}\,h(p_{0})v_{0},v_{0}\bigr)=\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\Big|_{\rm geod}h(p_{0})$
@@
Similarly, as the Euclidean case, $\alpha$-strongly geodesic convexity implies the gradient dominance condition:
(37) $g_{p}\bigl(\text{grad}\,h(p),\text{grad}\,h(p)\bigr)\geq 2\alpha\bigl(h(p)-h(p^{*})\bigr),$
where $p^{*}$ is the global minimizer of $h$. Again, this inequality leads to the exponential convergence of gradient flows of $h$ on the manifold $\mathcal{M}$.
3.2 Hessian operator in the Fisher-Rao geometry
@@
(38) $g_{\rho}^{\mathrm{FR}}(\sigma_{1},\sigma_{2})=\int\frac{\sigma_{1}\sigma_{2}}{\rho}\mathrm{d}\theta,\text{ for }\sigma_{1},\sigma_{2}\in T_{\rho}\mathcal{P}.$
e Fisher-Rao metric can also be viewed as an infinite dimensional Euclidean metric restricted to the unit ball $\{\psi:\int\psi^{2}=1\}$, if we apply the transformation $\psi=\sqrt{\rho}$; see, e.g., [halder2018gradient]. We note that one could derive the formulas of the geodesic convexity conditions and the relevant functional inequalities using the transformed coordinate $\psi$ by exploiting the Euclidean structure. In this paper, we choose to work with the original coordinates, as they eventually lead to equivalent formulas. Determining for which functions $f$ these conditions hold is a separate question independent from the coordinates chosen and is the one we primarily address in this work.
The geodesic distance between $\rho_{0},\rho_{1}$, referred to as the Fisher-Rao distance, satisfies
(39a) $\mathcal{D}_{\rm FR}^{2}(\rho_{0},\rho_{1})$ $=\inf_{\rho_{t},\sigma_{t}}\Bigl\{\int_{0}^{1}\int\frac{|\sigma_{t}|^{2}}{\rho_{t}}\mathrm{d}\theta\mathrm{d}t:\partial_{t}\rho_{t}=\sigma_{t}\Bigr\}$
(39b) $=\inf_{\rho_{t},\psi_{t}}\Bigl\{\int_{0}^{1}\int\rho_{t}\bigl(\psi_{t}-\mathbb{E}_{\rho_{t}}[\psi_{t}]\bigr)^{2}\mathrm{d}\theta\mathrm{d}t:\partial_{t}\rho_{t}=\rho_{t}(\psi_{t}-\mathbb{E}_{\rho_{t}}[\psi_{t}])\Bigr\}.$
The Fisher-Rao geodesic distance attains an explicit formula, equivalent to (up to a constant scaling) the spherical Hellinger distance [halder2018gradient, laschos2019geometric, lu2022birth]: $\mathcal{D}_{\rm FR}^{2}(\rho_{0},\rho_{1})=4\operatorname{arccos}^{2}\left(\int\sqrt{\rho_{0}}\sqrt{\rho_{1}}{\rm d}\theta\right)$. In view of this relation, Fisher-Rao gradient flows are sometimes referred to as spherical Hellinger gradient flows in the literature [lindsey2022ensemble, lu2022birth]. If we do not restrict the distributions $\rho_{t}$ to be on the probability space and we allow them to have any positive mass, then by using the relation $\frac{|\sigma_{t}|^{2}}{\rho_{t}}=\frac{|\partial_{t}{\rho}_{t}|^{2}}{\rho_{t}}=4\Bigl|\frac{\mathrm{d}}{\mathrm{d}t}\sqrt{\rho_{t}}\Bigr|^{2}$ and the Cauchy-Schwarz inequality, we will get that the optimal objective value in (39) is $4\int|\sqrt{\rho_{0}}-\sqrt{\rho_{1}}|^{2}\mathrm{d}\theta.$ This is (up to a constant scaling) the Hellinger distance [gibbs2002choosing]. For a more detailed discussion of the origin and historical development of the terms Fisher–Rao and Hellinger, we refer to [mielke2025hellinger, Remark 2.2].
Using a formal Lagrangian duality calculation, we can derive the optimality condition of the variational problem (39) as a coupled system of differential equations (see also [wang2020information]):
(40) $\partial_{t}\rho_{t}=\rho_{t}\bigl(\psi_{t}-\mathbb{E}_{\rho_{t}}[\psi_{t}]\bigr),\qquad\partial_{t}\psi_{t}=-\frac{1}{2}(\psi_{t}-\mathbb{E}_{\rho_{t}}[\psi_{t}])^{2},$
@@
(41a) $g_{\rho}^{\rm FR}(\textrm{grad}_{\rm FR}\mathcal{E}(\rho),\sigma)=$ $\frac{\mathrm{d}}{\mathrm{d}t}\Big|_{\rm geod}\mathcal{E}(\rho)=\int\frac{\delta\mathcal{E}}{\delta\rho}\rho\bigl(\psi-\mathbb{E}_{\rho}[\psi]\bigr)\mathrm{d}\theta,$
$g_{\rho}^{\rm FR}(\textrm{Hess}_{\rm FR}\mathcal{E}(\rho)\sigma,\sigma)=$ $\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\Big|_{\rm geod}\mathcal{E}(\rho)=\frac{\mathrm{d}}{\mathrm{d}t}\Big|_{\rm geod}\int\frac{\delta\mathcal{E}}{\delta\rho}\rho\bigl(\psi-\mathbb{E}_{\rho}[\psi]\bigr)\mathrm{d}\theta$
(41b) $=$ $\int\frac{\delta^{2}\mathcal{E}}{\delta\rho^{2}}\rho^{2}\bigl(\psi-\mathbb{E}_{\rho}[\psi]\bigr)^{2}+\frac{1}{2}\rho\bigl(\psi-\mathbb{E}_{\rho}[\psi]\bigr)^{2}\Bigl(\frac{\delta\mathcal{E}}{\delta\rho}-\mathbb{E}_{\rho}[\frac{\delta\mathcal{E}}{\delta\rho}]\Bigr)\mathrm{d}\theta.$
Here, we repeatedly used the derivatives of $\rho$ and $\psi$ along constant-speed geodesics in Eq. 40.
3.3 Geodesic convexity in the Fisher-Rao geometry
For the $f$-divergences (23), the corresponding gradient and Hessian operators are
(42a) $g_{\rho}^{\rm FR}\bigl(\textrm{grad}_{\rm FR}D_{f}[\rho\|{\rho^{*}}],\sigma\bigr)=$ $\int f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)\rho\bigl(\psi-\mathbb{E}_{\rho}[\psi]\bigr)\mathrm{d}\theta,$
(42b) $g_{\rho}^{\rm FR}\bigl(\textrm{Hess}_{\rm FR}D_{f}[\rho\|{\rho^{*}}]\sigma,\sigma\bigr)=$ $\int f^{\prime\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)\frac{\rho^{2}\bigl(\psi-\mathbb{E}_{\rho}[\psi]\bigr)^{2}}{{\rho^{*}}}$
$+$ $\frac{1}{2}\rho\bigl(\psi-\mathbb{E}_{\rho}[\psi]\bigr)^{2}\Bigl(f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)-\mathbb{E}_{\rho}[f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)]\Bigr)\mathrm{d}\theta,$
where $\sigma=\rho\bigl(\psi-\mathbb{E}_{\rho}[\psi]\bigr)$.
@@ 3.3 Geodesic convexity in the Fisher-Rao geometry
note that small neighborhood in Theorem 3.1 can also be defined using the Fisher-Rao distance $\mathcal{D}_{\rm FR}^{2}(\rho,\rho^{*})$. This is because we have the inequality $\mathcal{D}_{\rm FR}^{2}(\rho,\rho^{*})\leq 4\mathrm{KL}[\rho\|\rho^{*}]$. Indeed,
$\mathcal{D}^{2}_{\rm FR}(\rho,\rho^{*})$ $=4\arccos^{2}\Bigl(\int\sqrt{\rho\rho^{*}}d\theta\Bigr)\leq 4\Bigl[-2\log\int\sqrt{\rho\rho^{*}}d\theta\Bigr]\leq 4\rm{KL}[\rho\|\rho^{*}],$
here the first inequality follows from $\arccos^{2}(x)\leq-2\log x\,\forall\,0\leq x\leq 1$ with $x=\int\sqrt{\rho\rho^{*}}{\rm d}\theta$. This can be shown by defining $g(t)=-2\log\cos t-t^{2}$ for $t=\arccos(x)\in[0,\frac{\pi}{2}]$, and noting that $g^{\prime}(t)=2\tan t-2t\geq 0$, which implies $g(t)\geq g(0)=0$. The second inequality follows from the monotonicity of the Rényi divergence
$D_{\alpha}[\rho\|\rho^{*}]=\frac{1}{\alpha-1}\log\int\rho^{\alpha}{\rho^{*}}^{1-\alpha}{\rm d}\theta,$
which is nondecreasing in $\alpha$ [van2010renyi]. Hence $2\log\int\sqrt{\rho\rho^{*}}{\rm d}\theta=D_{\frac{1}{2}}[\rho\|\rho^{*}]\leq D_{1}[\rho\|\rho^{*}]=\rm{KL}[\rho\|\rho^{*}].$
Proof 3.2.
The KL-divergence corresponds to the $f$-divergence with $f(x)=x\log x$. In such case, the Hessian formula (42b) takes the form
(43) $\begin{split}g_{\rho}^{\rm FR}\bigl(\textrm{Hess}_{\rm FR}{\rm KL}[\rho\|{\rho^{*}}]\sigma,\sigma\bigr)&=\int\rho\bigl(\psi-\mathbb{E}_{\rho}[\psi]\bigr)^{2}\\ &+\frac{1}{2}\rho\bigl(\psi-\mathbb{E}_{\rho}[\psi]\bigr)^{2}\Bigl(\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr)\mathrm{d}\theta.\end{split}$
The positive semi-definiteness of the Hessian depends on the sign of $1+\frac{1}{2}\bigl(\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\bigr)$. For instance, the KL divergence is geodesically convex when this term is non-negative. A simple sufficient condition is $1/e\leq{\rho}/{{\rho^{*}}}\leq e$.
However, in general, this term does not have a definite sign, and the KL divergence will not be geodesically convex. To illustrate this, let us consider a specific case where $\psi=\log\frac{\rho}{{\rho^{*}}}$. In this scenario, (43) becomes
(44) $\mathcal{H}(\rho,{\rho^{*}})=\int\rho\Bigl(\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr)^{2}+\frac{1}{2}\rho\Bigl(\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr)^{3}\mathrm{d}\theta.$
We will now present two examples to demonstrate why the above term is not generally non-negative.
We begin by considering a simple pedagogical counterexample where ${\rho^{*}}(\theta)=\mathcal{N}(\theta;0,1)$ represents a 1D Gaussian density. We then construct $\rho(\theta)=\mathcal{N}(\theta;\mu,\sigma^{2})$ as another Gaussian distribution. The following calculations can be made:
(45) $\begin{split}&\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]=\frac{\sigma^{2}}{2}+\frac{\mu^{2}}{2}-\log\sigma-\frac{1}{2},\\ &\int\rho\Bigl(\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr)^{2}\mathrm{d}\theta=\frac{1}{2}(\sigma^{2}-1)^{2}+\sigma^{2}\mu^{2},\\ &\int\rho\Bigl(\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr)^{3}\mathrm{d}\theta=(\sigma^{2}-1)^{3}+3(\sigma^{2}-1)\sigma^{2}\mu^{2}.\end{split}$
Plugging (45) into (44), we get
@@
We can choose $\sigma^{2}<\frac{1}{3}$ and $\mu^{2}>\frac{(\sigma^{2}-1)^{2}}{1-3\sigma^{2}}$, which leads to $\mathcal{H}(\rho,{\rho^{*}})<0$.
Next, we present a general example beyond the Gaussian case to demonstrate the theorem. We will first construct a counterexample where the density is not continuous. We then mollify this example to obtain a smooth density.
@@
(46) $\rho_{r}:=\int_{B_{r}}{\rho^{*}}(\theta)\mathrm{d}\theta,$
@@
(47) $\rho(\theta)={\rho^{*}}(\theta)(x_{1}\mathbbm{1}_{B^{c}_{r}}(\theta)+x_{2}\mathbbm{1}_{B_{r}}(\theta))$
where $\mathbbm{1}_{A}$ is the indicator function of a set $A$ and $B_{r}$ denotes the ball with radius $r$. For $\rho(\theta)$ to be a probability density, it is necessary that
(48) $x_{1}\Bigl(1-\int_{B_{r}}{\rho^{*}}(\theta)\mathrm{d}\theta\Bigr)+x_{2}\int_{B_{r}}{\rho^{*}}(\theta)\mathrm{d}\theta=x_{1}(1-\rho_{r})+x_{2}\rho_{r}=1.$
In fact, for any choice of $x_{1}>1>x_{2}>0$, we can take some $r\in(0,+\infty)$ such that $\rho_{r}=\frac{x_{1}-1}{x_{1}-x_{2}}\in(0,1)$ which leads to (48).
[…]
@@
(49) ${\rm KL}[\rho\|{\rho^{*}}]\leq$ $\frac{x_{1}-x_{1}x_{2}}{x_{1}-x_{2}}\log x_{1}\leq\log x_{1}=\epsilon,$
$\mathcal{H}(\rho,{\rho^{*}})=$ $\rho_{r}(1-\rho_{r})x_{1}x_{2}\Bigl(\log\frac{x_{1}}{x_{2}}\Bigr)^{2}\Bigl(1-\frac{x_{1}+x_{2}-2x_{1}x_{2}}{2(x_{1}-x_{2})}(M+\epsilon)\Bigr)$